OpenStudy (anonymous):

I don't understand how to do this Direct Variation problem.. Please Help!!!! Medal & Fan!

OpenStudy (anonymous):

http://prntscr.com/5glj1g

OpenStudy (danjs):

Perimeter = 4*(side length)

OpenStudy (danjs):

the perimeter varies directly with the side length following a form y = k*x

OpenStudy (anonymous):

Mhmm?

OpenStudy (danjs):

y = 4*x

OpenStudy (anonymous):

OHHH

OpenStudy (danjs):

you see ?

OpenStudy (anonymous):

Thank you i got it correct.

OpenStudy (anonymous):

Yes.